# Optimization of deterministic photonic graph state generation via local operations

### v2

@@ Abstract
Introduction.—Graph states are a family of multi-qubit states where the entanglement structure between qubit nodes is characterized by the edges of a graph [1, 2]. Photonic graph states, with photons as graph nodes, are essential resources for the realization of measurement-based [3, 4] and fusion-based [5] fault-tolerant quantum computing as well as various quantum communication protocols [6, 7, 8, 9], including quantum repeaters. Photons are unique among qubit architectures as they are mobile qubits, a feature necessary for many of the mentioned applications, and are relatively inexpensive to create and manipulate, with long coherence times.
@@
Figure 1: Schematic of the circuit used for the sequential generation of photonic graph states. The main components are (i) single qubit gates on photons (yellow), (ii) emitter blocks representing arbitrary operations on the set of emitter qubits between emission events (green) , and (iii) emission events (blue) corresponding to a two-qubit emitter-photon gate.
@@
Devising the necessary operations between emission events for generating a target state is not trivial. As qubit connectivity is heavily constrained, this task is fundamentally different from a general quantum circuit synthesis problem of establishing a desired quantum channel [21, 22]. To date, only a limited number of resource-conscious generation protocols have been proposed, with the majority designed for specific categories of graphs [16, 23, 24, 25] and only a few for arbitrary states [26, 27, 28], with the optimization efforts primarily concerned with minimizing the number of emitters. Although the emitter count is a significant factor, there are other important metrics that can be considered in the circuit design, such as the required emitter coherence times and the number of two-qubit gates.
@@
In this paper, we propose a method to reduce the generation costs of photonic graph states, by leveraging the concept of local Clifford (LC) equivalency of graphs [29]. While LC equivalency has been previously employed to optimize the preparation of graph states [22, 30, 31], this is the first time that it is applied to the case of deterministic photonic-state generation, differing notably from non-photonic cases. The method works through identifying a set of alternative intermediate graphs which can be converted to the final target state with minimal overhead with respect to the photonic platform constraints, providing a selection of generation routes to choose from. While deterministic generation protocols such as Ref. [27] use Clifford circuits to generate a single fixed graph (the target), our approach differs by aiming to identify the best state to be generated, among all graphs in the local Clifford equivalency class. Any deterministic graph-to-circuit map can later be used to obtain the generation circuit. Moreover, we adopt a novel approach by employing graph theoretical metrics to correlate the structural characteristics of a graph with the cost parameters of its generating circuit, such as circuit size and depth. This is used to determine which intermediate graphs are more favorable for the process, filtered solely based on their shape. The key insight of this work is leveraging the size and correlations of the LC equivalency class, leading to efficient and substantial reductions in generation costs. This forms the basis of a versatile optimization framework for photonic state generation.
@@
A primary result is the identification of the optimized elementary circuit blocks for generating arbitrarily large repeater graph states (RGS) [9], achieving a 50% reduction in required two-qubit gates between emitters. This reduction promises a potential orders-of-magnitude increase in the success rate of quantum repeater protocols. Moreover, we present for the first time a polynomial-time algorithm to generate the full local equivalency class of an RGS. We elaborate on the wide-range applicability of the new method by showcasing an efficient correlation-oriented optimization within the local equivalency classes of random dense graphs. We achieve average reductions of 65% and 46% in unitary gate count and ‘emitter depth’ (defined below), respectively, in generation circuits of 15-node graphs, and observe similar advantages as the graph size increases.
@@
Background and preliminaries.—To introduce an unambiguous definition for the generation cost and the resources to optimize for, we consider the quantum circuit framework in Fig. 1. The circuit models emitter qubits as entanglement mediators to produce a desired photonic state, taking a sequential block form with each block indicated by a photon emission event. Between successive emissions, arbitrary operations are allowed on the emitter space, but only single-qubit unitaries can be applied on the emitted photons. Such a configuration, with enough emitters, has been proved to be sufficient for on-demand generation of arbitrary graph states [15]. To further simplify the cost analysis, the domain of operations considered in each step can be restricted to Clifford group unitaries and measurements in the computational basis, as they are proven to be universal for the preparation of graph states while being efficient to simulate classically [10, 32].
@@
It should be noted that although the photonic gates can be applied in a sequential manner in each step, for each photonic register, one can combine all single qubit gates throughout the circuit into a single unitary to be applied at once. Therefore, the photonic part of the circuit has minimal contribution to its size. Besides, the photonic gates are usually realized through linear optical elements, e.g., waveplates, and thus are easier to implement compared to the gates applied on the emitter qubits. Consequently, the emitter part of the circuit plays the primary role in cost considerations.
@@
Notable circuit parameters to consider in a cost analysis are (i) the number of emitters, (ii) the number of gates of each type, and (iii) the circuit depth. In particular, using more emitters necessitates coherent control over a larger set of qubits, and a wider extent of emitter-emitter couplings, which is challenging in practice. This emitter coupling cost can be directly reflected in the number of two-qubit gates which, apart from being difficult to implement, are also a primary source of infidelity in the circuit. The depth of the circuit is associated with the maximum number of successive gates applied on the qubits and is also indicative of the time it takes to complete the protocol; two significant factors in circuit performance. Since an emitter can undergo a reset after each measurement, as seen in protocols outlined in Refs. [27, 24], and its coherence time can be considered to reset as well, a more physically relevant metric would be the maximum number of consecutive gates on an emitter in the interval between each reset to |$\ket{0}$ and a subsequent measurement. Here, we define this new metric as ‘emitter depth’, complementary to the regular circuit depth.
@@
In our analysis, while maintaining the minimum emitter count, the focus is on other experimentally significant metrics that have been mostly overlooked in the existing literature on deterministic photonic state generation, e.g., the number of emitter-emitter cnots, and the emitter depth. Nevertheless, the optimization process can be easily adapted to incorporate other customized cost functions.
@@
Optimization framework.—Our optimization method is based on two main notions: the local equivalency of states, and correlations between the structure of a graph, which is quantified by graph theoretical metrics, and the parameters of its generating circuit.
@@
Two quantum states are locally equivalent if they can be converted into each other by local operations only, and are termed as locally Clifford (LC) equivalent if the operations are limited to single-qubit Clifford unitaries. It has been shown [29, 33] that two graph states |$\ket{G_{1}}$ and |$\ket{G_{2}}$ are LC equivalent, if and only if their respective graphs can be mapped into each other via the successive application of a graph transformation rule known as local complementation (LC rule) [34]. When the LC rule is applied on node a in a graph $G$, the subgraph comprised of the neighbourhood $N_{a}$ of a is complemented while other edges in the graph remain intact (see Fig. 2).
@@
Figure 2: Example of graph transformation via LC rule application on the blue node. The intermediate stage depicts the edge removal (dashed) and creation (solid) processes.
@@
Checking for the local equivalency of two graphs and finding a sequence of LC operations connecting them can be efficiently performed [35]. Convertibility under local complementation forms a valid equivalence relation on the graphs and the equivalency class for any graph, also called an orbit, can be obtained by successive application of the LC rule on its nodes [29, 33, 36].
@@
To compile an exhaustive array of alternative target graphs, the LC orbit is first identified via iterative application of the LC rule to graph nodes until no new graphs can be found. Subsequently, all possible emission orderings are considered for each orbit member. The set of alternative targets is then fed to a graph-to-circuit map in order to create a set of generating circuits, which are then analyzed to be ranked with respect to a chosen figure of merit. Our method takes the graph-to-circuit map as an input but operates at a higher level to expand the set of targets. While our proposed optimization scheme is mapping protocol-agnostic, for concreteness, we have first focused on the protocol in Ref. [27] in our analyses. This protocol guarantees the use of a minimal number of emitters for a given photon emission order, and runs in polynomial time, making it a suitable benchmark for evaluating cost reduction performance. Hereafter, for a given graph, comparisons are made between the original circuit obtained from [27]’s protocol and its optimized counterparts. Lastly, to showcase that the framework can accommodate any graph-to-circuit mapping, we also implement the protocol from Ref. [28] in the optimization, obtaining similar advantages to those achieved with Ref. [27].
Results.
@@
Figure 3: (a): Original RGS. (b, c): Partially transformed graphs corresponding to diagram (d)’s correlated part. (d): Number of cnots vs Number of edges for graphs in the orbit of a 100-qubit RGS. The red point is the original RGS. A linear trend is visible within a portion of the orbit. (e, f): The elementary circuit blocks to be applied on the 2 emitters required to generate graphs seen in (b) or (c), adding the photon pair $p_{2k}$ and $p_{2k+1}$ to the state under generation. Emissions are modeled by the cnots between an emitter and a photon. From (e) to (f), the rate of use of emitter-emitter cnots (shaded in circuit) changes from $1$ per photon to $1/2$ per photon emitted. (e) Generates a photon pair as in the repeated part of the original RGS structure, that is, a root and a leaf. (e) Generates a photon pair corresponding to a pair in the transformed part of the graph, indicated by the shaded area in (c). $H$, $P$, and $X$ gates are Hadamard, Phase, and Pauli $X$, and $U$ is the Clifford unitary $HPHPX$. The last operations in both blocks are a classically controlled measurement followed by a reset.
@@
In general, finding the exhaustive set of alternative graphs is not always feasible except for small-sized or highly symmetric cases. This is because the LC orbit can grow rapidly with the size of the graph. For instance, for linear and cycle graphs the scaling is exponential [34]. In fact, determining the size of the orbit of an arbitrary graph is a #P-complete problem [37]. Due to this complexity, previous investigations into the LC orbit have been limited to graphs with a maximum of 12 nodes [22, 36].
@@
We remark that for structured graphs, which can be parameterized by size, a full orbit analysis of smaller graphs can lead to identifying general patterns that are applicable to the larger graphs of the same type. One example of interest is the RGS, shown in Fig. 3(a), which enables all-photonic quantum repeater protocols and facilitates long-distance quantum communications [24, 38, 39]. Exploring the orbit of small RGSs shows that a near 50% reduction in the number of emitter-emitter cnots is obtainable (see Fig. 3). In Fig. 3(d), we plot the correlation diagram for the circuit cnot count as a function of the edge count of their respective graphs in the orbit for a 100-qubit RGS. The orbit can be divided into a linearly correlated region and an uncorrelated one (shown in the inset) with respect to the edge count. Figure 3 (b) and (c) correspond to the first two graphs belonging to the linearly correlated part. To obtain the graphs shown in (b) and (c), local complementations are applied on the nodes of the original RGS shown in (a) in order of $\{1,3,1\}$ and $\{1,3,5,7,1\}$, respectively. Here, the node labeling (starting from 00) also corresponds to the photon emission order. Similarly, to obtain the $j$-$th$ graph along the linearly correlated part ($j=0$ being the original RGS) we must apply the LC rule on the odd numbered nodes up to $14j-$ 1, followed by an additional LC on node $1$. Remarkably, breaking such an LC rule sequence at any node, results in a unique graph. By comparing the sequences leading to two consecutive graphs ($j$ and $j+1$) we observe that the $j+1$-$th$ graph has two extra nodes, $14j+$ 1 and $34j+$ 3, before the final LC on node 1 in its sequence compared to the $j$-$th$ one. Breaking the sequence at these two points correspond to two graphs belonging to the uncorrelated region of the orbit, constituting the inset of Fig. 3(d). Such generalization allows us to navigate the orbit to reach the graph with the minimum cnot count. This also provides, for the first time, an efficient polynomial-time algorithm to generate and access the entire orbit of an arbitrarily large RGS. (See supplementary section (A) [40].) It is noteworthy that the LC orbit identification task via brute-force search methods is not an efficient and scalable process in general. The generation circuit blocks in Fig. 3(d-e) demonstrate how such graph transformations can change the scaling of the required number of 2-qubit gates from $n$ to $n/2$ for an $n$-qubit RGS, as the emission of each photon pair belonging to the transformed part requires a single cnot instead of the original two. The final optimal graph shape can be achieved by sequentially applying the LC rule to all central nodes of an RGS for odd $n/2$ (i.e., $\{1,3,\ldots,n-1\}$). For even $n/2$, the last node in the sequence is replaced by its adjacent leaf, resulting in the set $\{1,3,\ldots,n-2\}$. The cnot count is as a result reduced from $n-3$ to $n/2-2$. This optimized scaling, identified systematically here, aligns with the scaling of the ad-hoc protocol described in Ref. [24], which was specifically designed for the optimal generation of RGS using a different approach and a different emission sequence. Furthermore, employing the mapping protocol of Ref. [28] instead of Ref. [27], results in similar scaling advantages in terms of circuit depth. (See Supplementary section (B) [40].)
The importance of our results lies in the ability to double the maximum size of the producible RGS, which is limited by the emitters’ finite coherence time, as the cnots between emitters are the primary time-consuming operations in such circuits [28]. The significance of this size increase is better reflected in the repeater protocol’s success probability in establishing an EPR pair between two distant parties, which is given by $P=(1-(1-P_{b})^{n/4})^{N}$ [9]. Here, $n$ is number of photons in RGS, $N$ is the number of RGS stations between the two parties, and $P_{b}$ is the probability of a successful entanglement swapping between two photons from two adjacent RGSs. Doubling the size from $n$ to $2n$, especially in the realistic long distance regime with $n$ being in the order of tens and $N$ in hundreds, results in several orders-of-magnitude increase in the overall success rate. (See Supplementary section (C) [40] for an example.)
To demonstrate the applicability of our method beyond structured graphs, we also investigated a set of graphs on the opposite end of the spectrum, i.e., random graphs [41]. Here, we limit the study to dense graphs with a 95% edge probability. Conjecturing a correlation between the generation cost and the shape of a graph, we chose a simple metric to parameterize the graph shape: the number of edges. This correlation can be directly confirmed for graphs of small size, e.g., 15 nodes, as the whole orbit is accessible. By choosing the graph with the least edge count in the orbit, we observed average reductions of $65\%\pm 11\%$ in the number of unitaries and $46\%\pm 18\%$ in emitter depth. One can efficiently test this correlation for larger graphs, as there are polynomial-time methods to find sparse graphs within the orbit of dense graphs (see supplementary section (D) [40] for one such strategy developed by the authors). Figure 4 shows the cost reduction achieved when extending the use of this strategy for graph of up to 65 nodes.
We also conducted an exhaustive analysis (see Supplementary Section E [40]) of graphs with up to 7 qubits [42], enumerating all the possible 45 entanglement classes as defined in Ref. [33] and comparing the generation circuits across $\sim 1.9\times 10^{6}$ labeled connected graphs [43]. All the presented analyses were conducted on a 3.2 GHz, 8-core, Apple M1 Pro chip with 16 GB of RAM, using our Python package “GraphiQ” [44].
Discussion.—We introduce an optimization framework tailored for the deterministic generation of photonic graph states, a task that poses significant theoretical and experimental challenges due to its inherent constraints. Our approach leverages the concept of local Clifford equivalency and local complementation to effectively identify alternative graphs in the same local equivalency class, that can serve as more resource-efficient paths for reaching a target state.
It is worth noting that the employed cost metrics can be tailored to meet the requirements of specific experimental platforms, e.g., as we have done with the ‘emitter depth’ to address the limitations in emitters’ coherence time. Additionally, the application can be easily extended to the probabilistic fusion-based generation protocols, as in Ref. [5], by facilitating the expansion of both the size scale and diversity of resource states used as building blocks.
Acknowledgement
Supplemental Materials
A: Repeater graph state analysis
The correlation analysis results in identifying interesting patterns in the case of repeater graph states. As evident in the Fig. 3(d) of main text, a subclass of the orbit shows almost perfect inverse linear correlation between the cnot and edge counts. By studying this subclass of the orbit, one can find the sequence of LC operations required so that the graph, and its corresponding circuit structure, moves one step toward the lower-cost region. Such an analysis gives rise to an efficient algorithm to not only generate the correlated subclass of the orbit, but also the full LC equivalency class for any RGS with arbitrary size. The procedure is outlined in detail in Algorithm 1 below.
@@ Acknowledgement
Figure S1: Seven first graphs in the orbit of a 9-qubit RGS. The numbers on arrows indicate on what node the local complementation should be applied to obtain the next graph. The observed pattern leads to the identification of an algorithm for a full orbit enumeration and being able to find elementary circuit blocks to realize each different shape. On top, from left to right, in each step a pair of arms from the RGS is folded into the central highly connected component. On bottom, from left to right, in each step one of the leaf nodes (blue nodes) gets connected to nodes 0 and 1, and the nodes 0 and 1 alternate between being connected and not connected. The new edges created in each step are represented with bold lines.
It is worth noting that determining the LC equivalency class of a graph is a non-trivial task; in fact, even determining the size of the orbit is a #P-complete problem [1]. In general as one needs to perform a depth or breadth-first search over sequences of local complementation operations on graph nodes, and check for isomorphisms of newly found graphs in each step which is itself an inefficient task. Access to the whole orbit enables one to confirm the results found through correlation analysis and moreover to perform exhaustive optimization with respect to different cost functions.
@@
Figure S2: The best alternative graph shapes in the orbit of an $n$-node RGS with odd (left) and even (right) number of arms ($n/2$= 9 and 8 respectively) in its original form. Larger cases can be obtained by adding pairs of nodes to the graphs such that the added two nodes are connected to each other and to the two central nodes.
@@
Figure S1 shows the pattern observed in the orbit and the required local complementation operations to navigate between them. The graphs on top correspond to the first cases in the linearly correlated part of Fig. 3(d). The graphs on bottom represent the uncorrelated part of Fig. 3(d) (the inset). By looking into the generating circuit for each case, one can find different elementary circuit blocks, as shown in the main article, that are to be utilized to add photons to different parts of the graph. Ultimately, the best (with respect to the use of cnot gates) alternative shape in the orbit and its respective generating circuit can be identified by continuing the trends seen in Fig. S1; the optimal case is depicted in Fig. S2 and Fig. S3.
Algorithm 1 RGS orbit enumeration

### v3

@@ Abstract
I Introduction
Graph states are a family of multi-qubit states where the entanglement structure between qubit nodes is characterized by the edges of a graph [1, 2]. Photonic graph states, with nodes representing photonic qubits, are essential resources for the realization of measurement-based [3, 4] and fusion-based [5] fault-tolerant quantum computing as well as various quantum communication protocols, including quantum repeaters [6, 7, 8, 9]. Photons are unique among qubit architectures as they are mobile qubits, a feature necessary for many of the mentioned applications, and are relatively inexpensive to create and manipulate, with long coherence times.
@@ I Introduction
Figure 1: Schematic of the circuit used for the sequential generation of photonic graph states. The main components in each emission block are (i) single qubit gates on photons, (ii) an emitter block representing arbitrary operations on the set of emitter qubits between photon emissions, and (iii) a two-qubit emitter-photon gate corresponding to a photon emission event.
@@
Devising the necessary operations between emission events for generating a target state is not trivial. As qubit connectivity is heavily constrained, this task differs fundamentally from general quantum circuit synthesis problems aimed at establishing a desired quantum channel [21, 22]. To date, only a limited number of resource-conscious generation protocols have been proposed, with the majority designed for specific categories of graphs [16, 23, 24, 25] and only a few for arbitrary states [26, 27, 28], with the optimization efforts concerned primarily with minimizing the number of emitters. Although the emitter count is a significant factor, there are other important metrics that can be considered in the circuit design, such as the required emitter coherence times and the number of two-qubit gates.
@@
In this paper, we propose a method to reduce the generation costs of photonic graph states, by leveraging the concept of local Clifford (LC) equivalency of graphs [29]. While LC equivalency has been previously employed to optimize the preparation of graph states [22, 30, 31], this is the first time that it is applied to the case of deterministic photonic-state generation, differing notably from non-photonic cases. The method works through identifying a set of alternative intermediate graphs which can be converted to the final target state with minimal overhead with respect to the photonic platform constraints, providing a selection of generation routes to choose from. While deterministic generation protocols such as Ref. [27] use Clifford circuits to generate a single fixed graph (the target), our approach differs by aiming to identify the best state to be generated, among all graphs in the local Clifford equivalence class. Any deterministic graph-to-circuit map can in principle be used to obtain the generation circuit for this process. Moreover, we adopt a novel approach by employing graph theoretical metrics to correlate the structural characteristics of a graph with the cost parameters of its generating circuit, such as circuit size and depth. This is used to determine which intermediate graphs are more favorable for the process, filtered solely based on their shape. The key insight of this work is leveraging the size and correlations of the LC equivalency class, leading to efficient and substantial reductions in generation costs. This lays the foundation for a versatile optimization framework for photonic state generation.
@@
A primary result is the identification of the optimized elementary circuit blocks for generating arbitrarily large repeater graph states (RGS) [9], achieving a 50% reduction in required two-qubit gates between emitters. This reduction promises a potential orders-of-magnitude increase in the success rate of quantum repeater protocols. Moreover, we present for the first time a polynomial-time algorithm to generate the full local equivalency class of an RGS. We elaborate on the wide-range applicability of the new method by showcasing an efficient correlation-oriented optimization within the local equivalency classes of random dense graphs. We achieve average reductions of 65% and 46% in unitary gate count and ‘emitter depth’ (defined in Section II), respectively, in generation circuits of 15-node graphs, and observe similar advantages as the graph size increases.
II Background and preliminaries
@@
To introduce an unambiguous definition for the generation cost and the resources to optimize for, we consider the quantum circuit framework in Fig. 1. The circuit models emitter qubits as entanglement mediators to produce a desired photonic state, taking a sequential block form with each block indicated by a photon emission event. Between successive emissions, arbitrary operations are allowed on the emitter space, but only single-qubit unitaries can be applied on the emitted photons. Such a configuration, with enough number of emitters, has been proved to be sufficient for on-demand generation of arbitrary graph states [15]. To further simplify the cost analysis, the domain of operations considered in each step can be restricted to Clifford group unitaries and measurements in the computational basis, as they are proven to be universal for the preparation of graph states while being efficient to simulate classically [10, 32].
@@
It should be noted that although the photonic gates can be applied in a sequential manner in each step, for each photonic register, one can combine all single qubit gates throughout the circuit into a single unitary to be applied at once. Therefore, the photonic part of the circuit has minimal contribution to its size. Additionally, the photonic gates are usually realized through linear optical elements, e.g., waveplates, and thus are easier to implement compared to the gates applied on the emitter qubits. Consequently, the emitter part of the circuit plays the primary role in cost considerations.
@@
Notable circuit parameters to consider in a cost analysis are (i) the number of emitters, (ii) the number of gates of each type, and (iii) the circuit depth. In particular, using more emitters necessitates coherent control over a larger set of qubits, and a wider extent of emitter-emitter couplings, which is challenging in practice. This emitter coupling cost can be directly reflected in the number of two-qubit gates which, apart from being difficult to implement, are also a primary source of infidelity in the circuit. The depth of the circuit is associated with the maximum number of successive gates applied on the qubits and also indicates the time required to complete the protocol; two significant factors in circuit performance. As seen in protocols outlined in Refs. [27, 24], an emitter qubit can undergo a reset after each measurement. Since the coherence time can be considered to reset as well in such cases, a more physically relevant metric would be the maximum number of consecutive gates applied on an emitter in the interval between each reset to |$\ket{0}$ and a subsequent measurement. Here, we define this new metric as ‘emitter depth’, complementary to the regular circuit depth.
@@
In our analysis, while maintaining the minimum emitter count, the focus is on other experimentally significant metrics that have mostly been overlooked in the existing literature on deterministic photonic state generation, e.g., the number of emitter-emitter cnots, and the emitter depth. Nevertheless, the optimization process can be easily adapted to incorporate other customized cost functions.
III Optimization framework
@@
Our optimization method is based on two main notions: the local equivalency of states, and correlations between the structure of a graph, which is quantified by graph theoretical metrics, and the parameters of its generating circuit.
@@
Two quantum states are locally equivalent if they can be converted into each other by local operations only, and are termed as locally Clifford (LC) equivalent if the operations are limited to single-qubit Clifford unitaries. It has been shown [29, 33] that two graph states |$\ket{G_{1}}$ and |$\ket{G_{2}}$ are LC equivalent, if and only if their respective graphs can be transformed into each other through the successive application of a graph transformation rule known as local complementation (LC rule) [34]. When the LC rule is applied on node a in a graph $G$, the subgraph comprised of the neighbourhood $N_{a}$ of a is complemented, while other edges in the graph remain intact (see Fig. 2). This operation corresponds to the local unitary [33]:
$U_{a}=\sqrt{-iX_{a}}\prod_{b\in N_{a}}\sqrt{iZ_{b}}$ (1)
in terms of Pauli operators $X_{a}$ and $Z_{b}$ acting on the nodes a and b.
@@
Figure 2: Example of graph transformation via local complementation on the black node in the initial graph on the left. The intermediate stage in the middle depicts the edge removal (dashed line) and edge creation (solid line) processes and the affected sub-graph denoted by the highlighted nodes.
@@
Checking for the local equivalency of two graphs and finding a sequence of LC operations connecting them can be efficiently performed [35]. Convertibility under local complementation forms a valid equivalence relation on the graphs. The equivalency class for any graph, also called an orbit, can be obtained by successive application of the LC rule on its nodes [29, 33, 36].
@@
To compile an exhaustive array of alternative target graphs, the LC orbit is first identified via iterative application of the LC rule to graph nodes until no new graphs can be found. Subsequently, all possible emission orderings are considered for each orbit member. The set of alternative targets is then fed to a graph-to-circuit map in order to create a set of generating circuits, which are subsequently analyzed to be ranked with respect to a chosen figure of merit. Our method takes the graph-to-circuit map as an input but operates at a higher level to expand the set of targets. While the proposed optimization scheme may utilize any mapping protocol, for concreteness, we have first focused on the protocol in Ref. [27] in our analyses. This protocol guarantees the use of a minimal number of emitters for a given photon emission order, and runs in polynomial time, making it a suitable benchmark for evaluating cost reduction performance. Hereafter, for a given graph, comparisons are made between the original circuit obtained from [27]’s protocol and its optimized counterparts. Lastly, to showcase that the framework can accommodate any graph-to-circuit mapping, we also implement the protocol from Ref. [28] in the optimization, obtaining similar advantages to those achieved with Ref. [27].
IV Results
@@
Figure 3: The original RGS is shown in (a), along with partially transformed graphs in (b) and (c), corresponding to the correlated part of the cnot behavior diagram shown in (d). In this diagram, the number of cnots is plotted against the number of edges for all the graphs in the orbit of a 100-qubit RGS, with the star-shaped marker indicating the original RGS. A linear correlation trend is visible within a portion of the orbit. The elementary circuit blocks in (e) and (f) are applied to the two emitters to generate the graphs shown in (b) or (c), adding the photon pair $p_{2k}$ and $p_{2k+1}$ to the state under generation. In these circuit blocks, emissions are modeled by the cnots between an emitter and a photon. From (e) to (f), the rate of use of emitter-emitter cnots (shaded in the circuit) changes from 1 per generated photon to 1/2. The circuit block in (e) generates a photon pair as in the repeated part of the original RGS structure, corresponding to a root and a leaf, while (f) generates a photon pair corresponding to the transformed part of the graph, indicated by the shaded pair of nodes in the graph (c). Gates $H$, $P$, and $X$ represent Hadamard, Phase, and Pauli-$X$, respectively, while $U$ is the Clifford unitary $HPHPX$. The final operations in both blocks are a measurement, a classically controlled gate based on the measurement result, followed by the emitter reset.
@@
In general, finding the exhaustive set of alternative graphs is not always feasible except for small-sized or highly symmetric cases. This is because the LC orbit can grow rapidly with the size of the graph. For instance, for linear and cycle graphs the scaling is exponential [34]. Due to this complexity, previous investigations into the LC orbit have been limited to graphs with a maximum of 12 nodes [22, 36]. However, identifying every member of the orbit is often unnecessary for gaining an advantage with respect to a given cost metric. In the following sections, we see cases where optimization can be achieved by recognizing certain patterns or correlations that allow us to find better performing graphs in a scalable and efficient manner without full orbit exploration.
IV.1 Repeater graph states
@@
For structured graphs that can be parameterized by size—i.e., graphs that preserve a characteristic node connectivity structure across different sizes—analyzing the full orbit of smaller instances can reveal general patterns applicable to larger graphs of the same type. The RGS, depicted in Fig. 3(a), is one such structured graph, facilitating all-photonic quantum repeater protocols and enabling long-distance quantum communication [24, 37, 38]. Exploring every member of the orbit of small RGSs indicates that the number of emitter-emitter cnot gates can be reduced by nearly 50%. This potential reduction can be seen in Fig. 3(d) where we plot a correlation diagram for the circuit cnot count as a function of the number of the edges in their respective graphs for the orbit of a 100-qubit RGS. We observe that the orbit can be divided into a linearly correlated region and an uncorrelated one (shown in the inset) with respect to the edge count. By studying the correlated portion of the orbit, one can find the sequence of LC operations required so that the graph, and the corresponding generating circuit, move one step toward the lower-cost region as indicated by the arrow in Fig. 3(d). Figure 3 (b) and (c) correspond to the first two graphs belonging to the linearly correlated part. To obtain the graphs shown in (b) and (c), local complementation operations (LC rule) are applied on the nodes of the original RGS shown in Fig. 3 (a) in order of $\{1,3,1\}$ and $\{1,3,5,7,1\}$, respectively. Here, the node labeling (starting from 00) corresponds to the photon emission order as well. Similarly, to obtain the $j$-$th$ graph along the linearly correlated part ($j=0$ being the original RGS) we must apply the LC rule on the odd numbered nodes up to $14j-$ 1, followed by an additional LC rule on node $1$. Remarkably, breaking such an LC rule sequence at any node, results in a unique graph in the local equivalency orbit. By comparing the sequences leading to two consecutive graphs ($j$ and $j+1$) we observe that the $j+1$-$th$ graph has two extra nodes, $14j+$ 1 and $34j+$ 3, before the final LC on node 1 in its sequence compared to the $j$-$th$ one. Stopping the transformation sequence at each of these two points results in two graphs belonging to the uncorrelated region of the orbit respectively, constituting the inset of Fig. 3(d).
@@
Figure 4: Seven first graphs in the orbit of a 18-qubit RGS. The numbers on arrows indicate on what node the local complementation is applied to obtain the next graph. The observed pattern leads to the identification of an algorithm for a full orbit enumeration and being able to find elementary circuit blocks to realize each different shape. On top, from left to right, in each step a pair of arms from the RGS is folded into the central highly connected component. On bottom, from left to right, in each step one of the leaf nodes (even numbered) gets connected to nodes 0 and 1, and the nodes 0 and 1 alternate between being connected and not connected. The new edges created in each step are represented with bold lines.
@@
Figure 5: The best alternative graph shapes in the orbit of an $n$-node RGS with odd (left) and even (right) number of arms ($n/2$= 9 and 8 respectively) in its original form. Larger cases can be obtained by adding pairs of nodes to the graphs such that the added two nodes are connected to each other and to the two central nodes.
@@
Figure 4 shows the pattern observed in the orbit and the required local complementation operations to navigate between them. The graphs in the top row correspond to the first few cases in the linearly correlated part of Fig. 3(d). The graphs on bottom represent the uncorrelated part of Fig. 3(d) (the inset).
{algorithm}
RGS orbit enumeration